OpenStudy (phebe):

What is the value of C so that y=x^2+15x+c is a perfect square trinomial? A-30 B-15/2 C- 15/4 D-225/4

OpenStudy (phebe):

@mjoxprm360

OpenStudy (anonymous):

what is half of \(15\) ?

OpenStudy (phebe):

5

OpenStudy (phebe):

5*3=15

OpenStudy (anonymous):

lets go slow what is HALF of 15, written as a fraction not "one third of fifteen"

OpenStudy (phebe):

um.. lol im not stupid but ok

OpenStudy (phebe):

7.5

OpenStudy (anonymous):

short answer!!!!! A

OpenStudy (phebe):

thanx @mjoxprm360

OpenStudy (anonymous):

i was looking for the answer "half of 15 is \(\frac{15}{2}\)" now to get your answer, you need to square it short answer is wrong, it is not A

OpenStudy (anonymous):

np

OpenStudy (phebe):

XD

OpenStudy (anonymous):

beware of answers given without explanation they are usually guesses and often wrong

OpenStudy (anonymous):

explain to me why its not.........................lol i forgot 2 numbers in my equation

OpenStudy (phebe):

foreal @satellite73 well not @mjoxprm360 he smart

OpenStudy (anonymous):

you need a method, not an answer any answer given without explanation is suspect in this case A is not the right answer, which is not surprising because there was no explanation of why it should be A, just the response "A"

OpenStudy (phebe):

foreal so thas wrong

OpenStudy (anonymous):

yes here ill put up explanation @satellite73

OpenStudy (phebe):

so wats the rit answer

OpenStudy (anonymous):

just a sec

OpenStudy (phebe):

ok

OpenStudy (anonymous):

lol phebe i put up wrong question XD XD XD XD XD XD XD

OpenStudy (anonymous):

half of \(15\) is \(\frac{15}{2}\) to get your answer, compute \[\left(\frac{15}{2}\right)^2\]